A certain job has been estimated to take 10 men 5days. On day 4, half of the job has been completed. How many additional men must be addd to the crew in order to complete the job on time?
Answer by palanisamy(496)

A certain job has been estimated to take 10 men 5days.
So the total number of mendays = 10*5 = 50
On day 4 half the job has been completed.
That means 25 mendays has been completed.
The remaining mendays = 50-25 = 25
In order to complete the work on the fifth day 25 men have to do work.
Therefore the number of additional men required = 25-10 = 15